### v2

@@ Abstract
Class-agnostic counting (CAC) aims to estimate the number of objects in images without being restricted to predefined categories. However, while current exemplar-based CAC methods offer flexibility at inference time, they still rely heavily on labeled data for training, which limits scalability and generalization to many downstream use cases. In this paper, we introduce CountingDINO, the first training-free exemplar-based CAC framework that exploits a fully unsupervised feature extractor. Specifically, our approach employs self-supervised vision-only backbones to extract object-aware features, and it eliminates the need for annotated data throughout the entire proposed pipeline. At inference time, we extract latent object prototypes via ROI-Align from DINO features and use them as convolutional kernels to generate similarity maps. These are then transformed into density maps through a simple yet effective normalization scheme. We evaluate our approach on the FSC-147 benchmark, where we consistently outperform a baseline based on an SOTA unsupervised object detector under the same label- and training-free setting. Additionally, we achieve competitive results – and in some cases surpass – training-free methods that rely on supervised backbones, non-training-free unsupervised methods, as well as several fully supervised SOTA approaches. This demonstrates that label- and training-free CAC can be both scalable and effective. Website: https://lorebianchi98.github.io/CountingDINO/.
@@ 1 Introduction
Class-agnostic counting (CAC) removes the need for class-specific retraining by allowing users to define target categories at inference [DBLP:conf/cvpr/RanjanSNH21]. This recent paradigm addresses the inherent limitations of class-specific counting approaches, which rely on predefined object types, such as vehicles [DBLP:conf/iccv/ZhangWCM17, DBLP:conf/iscc/AmatoCFG19], people [DBLP:conf/cvpr/LiuSF19, DBLP:journals/eswa/BenedettoCCFGA22], cells [DBLP:journals/mia/CiampiCTMLSAPG22, DBLP:conf/eccv/XueRHB16], or animals [DBLP:conf/eccv/ArtetaLZ16, DBLP:journals/cea/TianGCWLM19].
@@
However, CAC models still rely heavily on annotated training data. First, most existing approaches extract image features using backbones pre-trained on large-scale labeled datasets such as ImageNet [DBLP:conf/cvpr/HeZRS16]. Second, they require supervised training on annotated datasets, including thousands of dot annotations and bounding boxes, limiting scalability. Dot annotations, placed on object centroids, are used to generate density maps that serve as training targets for a regression network learning to map image features to object density estimates [DBLP:conf/nips/LempitskyZ10]. Bounding boxes, on the other hand, are used to localize a small set of visual prototypes, known as exemplars, which represent the object category to be counted [DBLP:conf/cvpr/RanjanSNH21]. Particularly, the requirement for thousands of dot annotations poses a major obstacle to creating large-scale datasets. As a result, only a few datasets are available, and even these often exhibit notable limitations [10944113].
@@
In this paper, we introduce CountingDINO, the first training-free exemplar-based CAC framework that eliminates the need for labeled data at any stage. Our image feature extraction backbone builds on DINO [caron2021emerging, oquab2023dinov2, darcet2024vision], which is trained in a self-supervised fashion. At inference time, we extract exemplar features by applying ROI-Align [8372616] to user-provided bounding boxes using the same DINO-computed image feature maps. These exemplar features are then used as depthwise convolutional kernels over the image features, generating similarity maps that highlight regions matching the exemplars. We average these similarity maps across exemplars to produce a global, informative response, which is then converted into a density map through a simple yet effective normalization scheme. Final object counts are computed by integrating the density map. Furthermore, to address the limitations of the spatial resolution of DINO with small objects, we partition the image into non-overlapping quadrants, apply feature extraction independently, and aggregate the resulting maps. We validate our method on the gold-standard CAC benchmark FSC-147 [DBLP:conf/cvpr/RanjanSNH21], consistently outperforming a baseline based on an SOTA unsupervised object detector under the same label- and training-free setting. We also compare it to training-free methods relying on supervised backbones, non-training-free unsupervised methods, and fully supervised approaches, demonstrating that our method remains competitive despite not using supervision or training at any stage.
@@
  • We assess our approach on the FSC-147 benchmark. CountingDINO outperforms a baseline under the same label- and training-free setting and achieves results that are comparable to – and in some cases surpass – recent training-free methods with supervised backbones, unsupervised approaches (that are not training-free), as well as fully supervised approaches.
@@ 2.1 Exemplar-based Class-agnostic Counting
Counting category-specific objects is a longstanding task in computer vision with broad real-world applications – e.g., counting of people [DBLP:conf/cvpr/LiuSF19, DBLP:journals/eswa/BenedettoCCFGA22], vehicles [DBLP:conf/iccv/ZhangWCM17, DBLP:conf/iscc/AmatoCFG19], insects [DBLP:journals/ecoi/CiampiZICBFAC23, DBLP:journals/cea/Bereciartua-Perez22], or biological structures [DBLP:journals/mia/CiampiCTMLSAPG22, DBLP:conf/eccv/XueRHB16]. Among existing counting approaches, density map regression – estimating counts via feature-to-density mapping – has proven effective in crowded scenes [DBLP:conf/eccv/ArtetaLZ16, DBLP:conf/nips/LempitskyZ10], outperforming detection-based techniques [DBLP:journals/cea/TianGCWLM19, DBLP:conf/iscc/AmatoCFG19]. However, their reliance on category-specific annotations and training limits scalability and generalization.
Class-agnostic counting (CAC) has recently shifted the object counting paradigm toward open-world contexts, enabling models to handle arbitrary categories unseen during training. In this setting, users are no longer restricted to predefined categories and can instead specify novel object classes at inference time by providing visual exemplars – typically in the form of three bounding boxes enclosing visual prototypes within the same input image [DBLP:conf/cvpr/RanjanSNH21]. The seminal FamNet [DBLP:conf/cvpr/RanjanSNH21] combines multi-scale feature extraction with exemplar-image correlation via inner product operations for density map prediction. RCAC [DBLP:conf/eccv/GongZ0DS22] follows a similar architecture but introduces a feature augmentation module that synthesizes exemplars with varied colors, shapes, and scales to enhance diversity. BMNet [DBLP:conf/cvpr/Shi0FL022] highlights the limitations of inner product-based correlation and instead proposes a learnable bilinear similarity loss, inspired by metric learning, for better matching. LOCA [DBLP:conf/iccv/EukicLZK23] introduces an object prototype extraction module that iteratively refines exemplar features via cross-attention, while PseCo [DBLP:conf/cvpr/HuangD0ZS24] relies on the popular SAM for instance segmentation [DBLP:conf/iccv/KirillovMRMRGXW23]. CACViT [DBLP:conf/aaai/WangX0024] leverages a single pre-trained Vision Transformer, using its attention mechanism for both feature extraction and matching. DAVE [DBLP:conf/cvpr/PelhanLZK24] proposes a two-stage detect-and-verify framework: first identifying candidate boxes via density maps and then verifying them using exemplar-based clustering. In contrast to these supervised methods, UnCounTR [knobel2024uncountr] is trained without manual annotations by using self-collages generated from unsupervised data and DINO features. Instead, TFPOC [DBLP:conf/wacv/Shi0Z24] and OmniCount [DBLP:journals/corr/abs-2403-05435] propose training-free pipelines using SAM without point-level supervision — though SAM is label-trained, making the approach not fully label-free. A recent survey of CAC methods is available in [DBLP:journals/corr/abs-2501-19184].
@@ 2.2 Unsupervised Vision Backbones
Recent advances in unsupervised, especially self-supervised, learning have produced models with highly effective representations for classification and localization. The DINO family [caron2021emerging, oquab2023dinov2, darcet2024vision], based on vision transformers [dosovitskiy2021an] and trained via self-distillation, demonstrates that patch-level features learned without labels can capture rich semantic information, outperforming CLIP-based models, which rely on weak supervision from image-text pairs. Alongside DINO, masked autoencoders (MAE) [he2022masked] offer a self-supervised approach by training transformers to reconstruct masked patches, fostering global visual understanding.
@@ 3 Method
Given an image and a small set of exemplar bounding boxes from the same image – each containing a visual prototype of the object we want to count – we aim to estimate the total number of instances of that object class present in the image, without relying on human-labeled data at any stage. To this end, we leverage the object understanding capabilities of DINO [caron2021emerging, oquab2023dinov2, darcet2024vision], a self-supervised vision-only backbone, and proceeding as described in the following sections. We also report a detailed schema in Fig. 1.
@@ 3.1 DINO-based Feature Extraction
To refine the exemplar representation, we apply a soft spatial prior in the form of an elliptical weighting mask centered within the bounding box. This mask attenuates peripheral regions and emphasizes the center under the assumption that exemplar objects are typically centered within their boxes. Formally, we define a mask $\mathbf{M}^{\mathbf{\tilde{b}}}\in\mathbb{R}^{w(\mathbf{\tilde{b}})\times h(\mathbf{\tilde{b}})}$, where each entry $m^{\mathbf{\tilde{b}}}_{ij}$ indicates the proportion of the corresponding feature cell in the bounding box $\mathbf{\tilde{b}}$ that lies within an ellipse centered in it. The impact of this spatial prior and the underlying assumption is further analyzed through ablation experiments in LABEL:ssec:ablation.
Figure 1: Overview of CountingDINO. Given an image $I$ and $N$ exemplar boxes, we extract features using the DINO-based visual backbone and apply each exemplar as a convolutional kernel over the image feature map to obtain similarity maps. These are aggregated, normalized into a density map using spatial priors, and thresholded before integration to produce the final count.
@@ 3.3 Density Map Normalization
The aggregated similarity map S well highlights regions likely to contain object instances. However, we seek a similarity response that integrates to 1 over each object instance, effectively transforming the similarity map into a density map from which the total count can be estimated via integration.
@@ 3.4 Thresholding density map contributions
However, notice that $u(\mathcal{B})$ assumes its minimum value when fed with the set =$\mathcal{B^{*}}=\{\operatorname*{arg\,max}_{\mathbf{b}\in\mathcal{B}}\text{area}(\mathbf{b})\}$ – i.e. when computed on the set composed solely of the largest exemplar bounding box. In cases different from this, we are also inevitably filtering out contributions from the other exemplars bounding boxes that we know are valid. For this reason, we set the threshold to be exactly the unit count computed on the largest region, i.e., $\tau=u(\mathcal{B}^{*})$.
@@ 3.5 Increasing Spatial Resolution
This modification integrates seamlessly into the pipeline, as all subsequent steps – ROI-Align, similarity computation, and normalization – remain unchanged. We found $k=2$ to work well in the presented domain, although we experiment with other values in LABEL:ssec:ablation.
@@ 4 Experimental Evaluation
4.1 Datasets and Metrics
We evaluate our method on the FSC-147 dataset [DBLP:conf/cvpr/RanjanSNH21], which consists of 6,135 images belonging to 147 object classes, each accompanied by three exemplar bounding boxes. We evaluate counting performance using two standard metrics: Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) [DBLP:conf/eccv/ArtetaLZ16, DBLP:conf/nips/LempitskyZ10, DBLP:conf/cvpr/RanjanSNH21], defined as $MAE=\frac{1}{n}\sum_{i=1}^{n}|c_{i}-\hat{c}_{i}|$ and $RMSE=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(c_{i}-\hat{c}_{i})^{2}}$, where $c_{i}$ and $\hat{c}_{i}$ are the ground truth and predicted counts for the $i$-th image.

### v3

@@ Abstract
Class-agnostic counting (CAC) aims to estimate the number of objects in images without being restricted to predefined categories. However, while current exemplar-based CAC methods offer flexibility at inference time, they still heavily rely on labeled data for training, which limits scalability and generalization to many downstream use cases. In this paper, we introduce CountingDINO, the first training-free exemplar-based CAC framework that exploits a fully unsupervised feature extractor. Specifically, our approach employs self-supervised vision-only backbones to extract object-aware features, and it eliminates the need for annotated data throughout the entire proposed pipeline. At inference time, we extract latent object prototypes via ROI-Align from DINO features and use them as convolutional kernels to generate similarity maps. These are then transformed into density maps through a simple yet effective normalization scheme. We evaluate our approach on the FSC-147 and CARPK benchmarks, where we consistently outperform a baseline based on an SOTA unsupervised object detector under the same label- and training-free setting. Additionally, we achieve competitive results – and in some cases surpass – training-free methods that rely on supervised backbones, non-training-free unsupervised methods, as well as several fully supervised SOTA approaches. This demonstrates that label- and training-free CAC can be both scalable and effective. Website: https://lorebianchi98.github.io/CountingDINO/.
@@ 1 Introduction
Class-agnostic counting (CAC) removes the need for class-specific retraining by allowing users to define target categories at inference [29]. This recent paradigm addresses the inherent limitations of class-specific counting approaches, which rely on predefined object types, such as vehicles [36, 1], people [24, 4], cells [7, 35], or animals [2, 32].
@@
However, CAC models still rely heavily on annotated training data. First, most existing approaches extract image features using backbones pre-trained on large-scale labeled datasets such as ImageNet [15]. Second, they require supervised training on annotated datasets, including thousands of dot annotations and bounding boxes, limiting scalability.
Figure 1: CountingDINO is a training-free and unsupervised class-agnostic counting method. The instance running with DINOv2 is able to outperform a training-free unsupervised detection-based baseline (CutLER), on both MAE and RMSE metrics (lower is better). Notably, it reaches comparable results (and even outperforms) previous non-training-free or supervised methods.
Dot annotations, placed on object centroids, are used to generate density maps that serve as training targets for a regression network learning to map image features to object density estimates [22]. Bounding boxes, on the other hand, are used to localize a small set of visual prototypes, known as exemplars, which represent the object category to be counted [29]. Particularly, the requirement for thousands of dot annotations poses a major obstacle to creating large-scale datasets. As a result, only a few datasets are available, and even these often exhibit notable limitations [10]. Many previous works tried to approach this level of unsupervision for the counting task. Nevertheless, in the end, they surrendered to performing supervised finetuning employing a variable amount of ground-truth density maps [23], or employing upstream vision backbones or segmentation models requiring costly annotations to be trained (e.g., SAM) [31].
@@
In this paper, we introduce CountingDINO, the first training-free exemplar-based CAC framework that eliminates the need for labeled data at any stage. CountingDINO shows that it is possible to effectively employ modern, powerful self-supervised vision features without any further downstream training or finetuning, or without relying on any kind of manually-labeled data. CountingDINO builds upon a smart and simple design, and shows that smart inference techniques can enable off-the-shelf self-supervised vision backbones to effectively count arbitrary objects in an image. Our image feature extraction backbone builds on DINO [6, 27, 11], which is trained in a self-supervised fashion. At inference time, we extract exemplar features by applying ROI-Align [16] to user-provided bounding boxes using the same DINO-computed image feature maps. These exemplar features are then used as depthwise convolutional kernels over the image features, generating similarity maps that highlight regions matching the exemplars. We average these similarity maps across exemplars to produce a global, informative response, which is then converted into a density map through a simple yet effective normalization scheme. Final object counts are computed by integrating the density map. Furthermore, to address the limitations of the spatial resolution of DINO with small objects, we partition the image into non-overlapping quadrants, apply feature extraction independently, and aggregate the resulting maps. We validate our method on the gold-standard CAC benchmarks FSC-147 [29] and CARPK [18], consistently outperforming a baseline based on an SOTA unsupervised object detector under the same label- and training-free setting. We also compare it to training-free methods relying on supervised backbones, non-training-free unsupervised methods, and fully supervised approaches, demonstrating that our method remains competitive despite not using supervision or training at any stage.
@@
  • We assess our approach on the FSC-147 and CARPK benchmarks. CountingDINO outperforms a baseline under the same label- and training-free setting and achieves results that are comparable to – and in some cases surpass – recent training-free methods with supervised backbones, unsupervised approaches (that are not training-free), as well as fully supervised approaches.
@@ 2.1 Exemplar-based Class-agnostic Counting
Counting category-specific objects is a longstanding task in computer vision with broad real-world applications – e.g., counting of people [24, 4], vehicles [36, 1], insects [8, 5], or biological structures [7, 35]. Among existing counting approaches, density map regression – estimating counts via feature-to-density mapping – has proven effective in crowded scenes [2, 22], outperforming detection-based techniques [32, 1]. However, their reliance on category-specific annotations and training limits scalability and generalization.
Class-agnostic counting (CAC) has recently shifted the object counting paradigm toward open-world contexts, enabling models to handle arbitrary categories unseen during training. In this setting, users are no longer restricted to predefined categories and can instead specify novel object classes at inference time by providing visual exemplars – typically in the form of three bounding boxes enclosing visual prototypes within the same input image [29]. The seminal FamNet [29] combines multi-scale feature extraction with exemplar-image correlation via inner product operations for density map prediction. RCAC [14] follows a similar architecture but introduces a feature augmentation module that synthesizes exemplars with varied colors, shapes, and scales to enhance diversity. BMNet [30] highlights the limitations of inner product-based correlation and instead proposes a learnable bilinear similarity loss, inspired by metric learning, for better matching. CounTR [23] proposes a method for counting elements in images of arbitrary semantic categories, by employing a two-stage training pipeline: a MAE-based self-supervised stage, which learns powerful visual features, and a fully-supervised final stage. LOCA [12] introduces an object prototype extraction module that iteratively refines exemplar features via cross-attention, while PseCo [19] relies on the popular SAM for instance segmentation [20]. CACViT [34] leverages a single pre-trained Vision Transformer, using its attention mechanism for both feature extraction and matching. DAVE [28] proposes a two-stage detect-and-verify framework: first identifying candidate boxes via density maps and then verifying them using exemplar-based clustering. In contrast to these supervised methods, UnCounTR [21] is trained without manual annotations by using self-collages generated from unsupervised data and DINO features. Instead, TFPOC [31] and OmniCount [25] propose training-free pipelines using SAM without point-level supervision — though SAM is label-trained, making the approach not fully label-free. A recent survey of CAC methods is available in [9].
@@ 2.2 Unsupervised Vision Backbones
Recent advances in unsupervised, especially self-supervised, learning have produced models with highly effective representations for classification and localization. The DINO family [6, 27, 11], based on vision transformers [13] and trained via self-distillation, demonstrates that patch-level features learned without labels can capture rich semantic information, outperforming CLIP-based models, which rely on weak supervision from image-text pairs. Alongside DINO, masked autoencoders (MAE) [17] offer a self-supervised approach by training transformers to reconstruct masked patches, fostering global visual understanding.
@@ 3 Method
Given an image and a small set of exemplar bounding boxes from the same image – each containing a visual prototype of the object we want to count – we aim to estimate the total number of instances of that object class present in the image, without relying on human-labeled data at any stage. We claim that, by leveraging the object understanding capabilities of DINO [6, 27, 11] – a family of self-supervised vision-only backbones exhibiting strong localized understanding capabilities – we can build a simple yet effective exemplar-based CAC model. While the core of the proposed approach is mainly constituted by the powerful DINO features, the proposed pipeline integrates simple yet clever downstream processing steps that can count objects given a few exemplars. We report a detailed schema in Fig. 2, and we describe each component in the following sections.
@@ 3.1 DINO-based Feature Extraction
To refine the exemplar representation, we apply a soft spatial prior in the form of an elliptical weighting mask centered within the bounding box. This mask attenuates peripheral regions and emphasizes the center under the assumption that exemplar objects are typically centered within their boxes. Formally, we define a mask $\mathbf{M}^{\mathbf{\tilde{b}}}\in\mathbb{R}^{w(\mathbf{\tilde{b}})\times h(\mathbf{\tilde{b}})}$, where each entry $m^{\mathbf{\tilde{b}}}_{ij}$ indicates the proportion of the corresponding feature cell in the bounding box $\mathbf{\tilde{b}}$ that lies within an ellipse centered in it. The impact of this spatial prior and the underlying assumption is further analyzed through ablation experiments in Section 4.5.
Figure 2: Overview of CountingDINO. Given an image $I$ and NN exemplar boxes, we extract features using the DINO-based visual backbone and apply each exemplar as a convolutional kernel over the image feature map to obtain similarity maps. These are aggregated, normalized into a density map using spatial priors, and thresholded before integration to produce the final count.
@@ 3.3 Density Map Normalization
The aggregated similarity map S well highlights regions likely to contain object instances. However, similarly to the final correction term applied in [23], we seek a similarity response that integrates to 1.0 over each object instance, effectively transforming the similarity map into a density map from which the total count can be estimated via integration.
@@ 3.4 Thresholding density map contributions
However, notice that $u(\mathcal{B})$ assumes its minimum value when fed with the set $\mathcal{B^{*}}=\{\operatorname*{arg\,max}_{\mathbf{b}\in\mathcal{B}}\text{area}(\mathbf{b})\}$ – i.e. when computed on the set composed solely of the largest exemplar bounding box. In cases different from this, we are also inevitably filtering out contributions from the other exemplars bounding boxes that we know are valid. For this reason, we set the threshold to be exactly the unit count computed on the largest region, i.e., $\tau=u(\mathcal{B}^{*})$.
@@ 3.5 Increasing Spatial Resolution
This modification integrates seamlessly into the pipeline, as all subsequent steps – ROI-Align, similarity computation, and normalization – remain unchanged. We found $k=2$ to work well in the presented domain, although we experimented with other values in Section 4.5.
@@ 4 Experimental Evaluation
4.1 Datasets
We evaluate our method on two widely-adopted datasets in the class-agnostic counting literature: FSC-147 [29] and CARPK [18].
FSC-147
The FSC-147 dataset [29] contains 6,135 images spanning 147 categories (e.g., plants, animals, vehicles). Each image includes three bounding boxes that serve as exemplars for a single object category. The number of objects per image varies significantly, from 7 to 3,731, with an average of 56. The data is split into training (3,659 images, 89 categories), validation (1,286 images, 29 categories), and test sets (1,190 images, 29 categories).
CARPK
The CARPK dataset [18] comprises 1,448 drone-view images of parking lots, designed for vehicle counting tasks. It was first employed in the context of CAC by [29], primarily to evaluate cross-dataset generalization when counting a specific object class – vehicles. The dataset includes approximately 90,000 annotated vehicles, each labeled with a bounding box. Density maps are generated by placing a point at the center of each bounding box, and a subset of these boxes is used as exemplars. The dataset is split into a training set of 989 images and a test set of 459 images.
4.2 Metrics
We evaluate counting performance using two standard metrics: Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) [2, 22, 29], defined as $MAE=\frac{1}{n}\sum_{i=1}^{n}|c_{i}-\hat{c}_{i}|$ and $RMSE=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(c_{i}-\hat{c}_{i})^{2}}$, where $c_{i}$ and $\hat{c}_{i}$ are the ground truth and predicted counts for the $i$-th image.